# GTPO and GRPO-S: Token and Sequence-Level Reward Shaping with Policy Entropy

### v5

@@ B.4 Gradient
Appendix C Group Relative Overlong Punishment: A Heuristic for Length Control
@@
Appendix D Additional Experimental Results and Analysis
D.1 Analysis of Training Dynamics: Entropy Rebound and Exploration
Figure 5 provides empirical evidence for the core mechanism of our proposed methods. The top row illustrates the ”entropy rebound” phenomenon. While all methods initially exhibit a decrease in policy entropy as they learn to exploit correct strategies, the DAPO baseline’s entropy continues to decline, indicating convergence to a narrow, deterministic policy—a behavior known as policy collapse. In contrast, both GTPO and GRPO-S show a distinct rebound in entropy. This is direct evidence that our entropy-weighted reward shaping successfully incentivizes the model to maintain exploration. By rewarding uncertainty on successful paths, our methods encourage the model to escape local optima and explore a more diverse set of reasoning pathways. This sustained exploration, as shown in the bottom row, directly results in longer, more detailed responses as the model attempts more thorough lines of reasoning. This establishes a clear causal link from our reward mechanism to superior problem-solving capabilities, as validated by the results in Table 1.
D.2 Analysis of Generation Characteristics: Response Length and Clipping
@@
D.3 Complete Reward Trajectories on Test Sets
@@
Figures 7 and 8 present the mean reward trajectories for Pass@4 and Pass@16, respectively. These plots complete the picture presented in Figure 3 of the main paper. The observed trends are highly consistent across all Pass@k metrics: GTPO and GRPO-S consistently achieve a higher final reward ceiling than the DAPO baseline, demonstrating the robustness of our performance gains.
D.4 Reward Trajectories on Training Sets
@@
Appendix E Qualitative Case Study
@@
This case study provides a compelling qualitative illustration of the practical difference between coarse-grained and fine-grained credit assignment. The DAPO model’s failure is characteristic of a system that has learned the general structure of a solution (the ”template”) but lacks logical rigor in its execution. Its coarse-grained reward signal (a single +1 or -1 for the entire sequence) is insufficient to penalize subtle but critical errors like the incorrect binary decomposition. The model can thus become overconfident in a flawed reasoning path. In contrast, the success of the GTPO and GRPO-S models highlights the benefit of an entropy-aware reward signal. Our methods are designed to penalize low-entropy (high-confidence) mistakes, which would directly discourage the kind of confident but incorrect decomposition made by the DAPO model. Simultaneously, by rewarding exploration in successful paths, our methods encourage a more careful and deliberate reasoning process, leading to the discovery of the correct, logically sound solution. This case demonstrates that our framework is key to moving LLMs beyond mere pattern imitation towards robust, verifiable reasoning.

### v6

@@ Appendix C Alternative Definitions via Geometric Mean
Appendix D Group Relative Overlong Punishment: A Heuristic for Length Control
@@
Appendix E Additional Experimental Results and Analysis
E.1 Analysis of Training Dynamics: Entropy Rebound and Exploration
Figure 6 provides empirical evidence for the core mechanism of our proposed methods. The top row illustrates the ”entropy rebound” phenomenon. While all methods initially exhibit a decrease in policy entropy as they learn to exploit correct strategies, the DAPO baseline’s entropy continues to decline, indicating convergence to a narrow, deterministic policy, often termed policy collapse. In contrast, both GTPO and GRPO-S show a distinct rebound in entropy. This is direct evidence that our entropy-weighted reward shaping successfully incentivizes the model to maintain exploration. By rewarding uncertainty on successful paths, our methods encourage the model to escape local optima and explore a more diverse set of reasoning pathways. This sustained exploration, as shown in the bottom row, directly results in longer, more detailed responses as the model attempts more thorough lines of reasoning. This establishes a clear causal link from our reward mechanism to superior problem-solving capabilities, as validated by the results in Table 1.
E.2 Analysis of Generation Characteristics: Response Length and Clipping
@@
E.3 Complete Reward Trajectories on Test Sets
@@
Figures 8 and 9 present the mean reward trajectories for Pass@4 and Pass@16, respectively. These plots complete the picture presented in Figure 3 of the main paper. The observed trends are highly consistent across all Pass@k metrics: GTPO and GRPO-S consistently achieve a higher final reward ceiling than the DAPO baseline, demonstrating the robustness of our performance gains.
E.4 Reward Trajectories on Training Sets
@@
Appendix F Qualitative Case Study
We present a qualitative case study to demonstrate the superiority of our GTPO and GRPO-S algorithms, formulated under the Dynamic Entropy Weighting framework, over DAPO. Below, we illustrate the actual problem-solving trajectories of these three algorithms on a mathematical problem requiring long-chain reasoning.
@@ Appendix F Qualitative Case Study
This case study provides a compelling qualitative illustration of the practical difference between coarse-grained and fine-grained credit assignment. The DAPO model’s failure is characteristic of a system that learns the general structure of a solution (the ”template”) but lacks logical rigor in its execution. Its coarse-grained reward signal (a single +1 or -1 for the entire sequence) is insufficient to penalize subtle but critical errors like the incorrect binary decomposition. The model can thus become overconfident in a flawed reasoning path. In contrast, the success of the GTPO and GRPO-S models highlights the benefit of an entropy-aware reward signal. Our methods are designed to penalize low-entropy (high-confidence) mistakes, which would directly discourage the kind of confident but incorrect decomposition made by the DAPO model. Simultaneously, by rewarding exploration in successful paths, our methods encourage a more careful and deliberate reasoning process, leading to the discovery of the correct, logically sound solution. This case demonstrates that our framework is key to moving LLMs beyond mere pattern imitation towards robust, verifiable reasoning.